A container with glass walls has oil inside. The total length,  of the containers is 14 cm. The walls of the container have a thickness () of 25 mm and a refractive index of 1.58. The container is surrounded by air (n = 1), and the oil inside the container has a refractive index of 1.42. The angle of incidence,  is 30º. Calculate the lateral shift, x, for this container. (5 marks)

1
Expert's answer
2022-01-14T09:48:52-0500

"x=(\frac {\sqrt{n_2^2-n_1^2}}nl-h)(1-\sin^2\alpha)=3.4~mm."